### v2

@@ A.1 Proof of Lemma 2
The upper bounds for $d\neq 2$ directly follow from Lemma 1, and it remains to prove the upper bound for $d=2$. We use $\mathbb{U}$ to denote uniform distribution. Recall that the Ajtai-Komlós-Tusnády Matching Theorem [AKT84] states that $\mathbb{E}_{S,R\sim{\mathbb{U}}^{n}}[\mathrm{OPT}(S,R)]=\Theta(\sqrt{n\log n})$ for $d=2$. We will generalize the proof of the upper bound in the Ajtai-Komlós-Tusnády Matching Theorem interpreted by [Tal92, Section 4.3].
We first set up some notations necessary for the proof. For all metric space $(X,\delta)$ and $X^{\prime}\subseteq X$, let $\Delta(X^{\prime})$ denote the diameter of $X^{\prime}$ with respect to $\delta$, i.e.,
$\Delta(X^{\prime})=\sup_{x,y\in X^{\prime}}\delta\left(x,y\right)\,.$ (1)
Given a set $X$, an admissible sequence of $X$ is an increasing sequence $({\mathcal{A}}_{t})_{t\geq 0}$ of partitions of $X$ where the cardinality of ${\mathcal{A}}_{t}$ satisfies $|{\mathcal{A}}_{0}|=1$ and $|{\mathcal{A}}_{t}|\leq 2^{2^{t}}$ for $t\geq 1$. Given an admissible sequence $({\mathcal{A}}_{t})_{t\geq 0}$ of $X$, we use $A_{t}(x)$ to denote the unique element of ${\mathcal{A}}_{t}$ that contains $x$ for all $t\geq 0$ and $x\in X$. For all metric space $(X,\delta)$ and $\alpha>0$, define
$\gamma_{\alpha}\left(X,\delta\right)$ $\triangleq\inf\sup_{x\in X}\sum_{t\geq 0}2^{t/\alpha}\Delta\left(A_{t}(x)\right)\,,$ (2)
where the infimum is taken over all admissible sequences of $X$.
Lemma 14.
For all $\alpha>0$, and metric spaces $(X,\delta_{1})$ and $(X,\delta_{2})$, if $\delta_{1}(x,y)\leq B\cdot\delta_{2}(x,y)$ for some $B>0$ and all $x,y\in X$, then $\gamma_{\alpha}\left(X,\delta_{1}\right)\leq B\cdot\gamma_{\alpha}\left(X,\delta_{2}\right)$.
Proof.
By (1) and (2),
$\gamma_{\alpha}\left(X,\delta_{1}\right)$ $=\inf\sup_{x\in X}\sum_{t\geq 0}2^{t/\alpha}\left(\sup_{y,z\in A_{t}(x)}\delta_{1}\left(y,z\right)\right)$
$\leq B\cdot\inf\sup_{x\in X}\sum_{t\geq 0}2^{t/\alpha}\left(\sup_{y,z\in A_{t}(x)}\delta_{2}\left(y,z\right)\right)=B\cdot\gamma_{\alpha}\left(X,\delta_{2}\right)\,,$
where the inequality holds by $\delta_{1}(x,y)\leq B\cdot\delta_{2}(x,y)$ for all $x,y\in X$. ∎
Now, we are ready to prove the upper bound for $d=2$. Let $X=[0,1]^{2}$, $\delta(x,y)=\left\|{x-y}\right\|_{2}$ for all $x,y\in X$, and $\mathbb{D}$ be a $\beta$-smooth distribution on $X$. Our goal is to prove that
$\mathrm{OPT}(n)=\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\leq O\left(\sqrt{n\log n}\right)\,.$
Let ${\mathcal{C}}$ denote the class of $1$-lipschitz functions on $(X,\delta)$, i.e., functions $f$ such that $|f(x)-f(y)|\leq\delta(x,y)$ for all $x,y\in X$. For all $f\in{\mathcal{C}}$ and $S=\{s_{1},\ldots,s_{n}\}\subseteq X$, define
$Z_{f}(S)\triangleq\sum_{i=1}^{n}\left(f(s_{i})-\mathbb{E}_{x\sim\mathbb{D}}\left[f(x)\right]\right)\,.$
By [Tal92, Proposition $4.2.3$], for all $S=\{s_{1},\ldots,s_{n}\},R=\{r_{1},\ldots,r_{n}\}\subseteq X$,
$\mathrm{OPT}(S,R)$ $=\sup_{f\in{\mathcal{C}}}\sum_{i=1}^{n}\left(f(s_{i})-f(r_{i})\right)$
$\leq\sup_{f\in{\mathcal{C}}}\sum_{i=1}^{n}\left(f(s_{i})-\mathbb{E}_{x\sim\mathbb{D}}\left[f(x)\right]\right)+\sup_{f\in{\mathcal{C}}}\sum_{i=1}^{n}\left(f(r_{i})-\mathbb{E}_{x\sim\mathbb{D}}\left[f(x)\right]\right)$
$=\sup_{f\in{\mathcal{C}}}|Z_{f}(S)|+\sup_{f\in{\mathcal{C}}}|Z_{f}(R)|\,.$ (3)
Analogous to [Tal92, Corollary $4.3.5$], for every $v>0$, we can show that
$\mathbb{P}_{S\sim\mathbb{D}^{n}}[\left|Z_{f}(S)\right|\geq v]\leq 2\exp\left(-\min\left(\frac{v^{2}}{4n\left\|{f}\right\|_{2}^{\mathbb{D}}},\frac{v}{4\left\|{f}\right\|_{\infty}^{\mathbb{D}}}\right)\right)\,,$ (4)
where $\left\|{f}\right\|_{p}^{\mathbb{D}}$ denotes the norm of $f$ in $L_{p}$ with respect to the measure $\mu_{\mathbb{D}}$ induced by the distribution $\mathbb{D}$. The proof of (4) is identical to the proof of [Tal92, Corollary $4.3.5$] by applying Bernstein’s inequality and hence omitted here. Then, analogous to [Tal92, Theorem $4.3.6$], for every class ${\mathcal{F}}$ of functions on $[0,1]^{2}$ with $0\in{\mathcal{F}}$, we can show that
$\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\sup_{f\in{\mathcal{F}}}\left|Z_{f}(S)\right|\right]\leq O\left(\sqrt{N}\gamma_{2}\left({\mathcal{F}},d_{2}^{\mathbb{D}}\right)+\gamma_{1}\left({\mathcal{F}},d_{\infty}^{\mathbb{D}}\right)\right)\,,$ (5)
where $d_{2}^{\mathbb{D}}$ and $d_{\infty}^{\mathbb{D}}$ are the distances induced on ${\mathcal{F}}$ by the norms of $L_{2}$ and $L_{\infty}$ respectively, with respect to the measure $\mu_{\mathbb{D}}$ induced by $\mathbb{D}$ on $[0,1]^{2}$, and $\gamma_{\alpha}$ is defined in (2) for $\alpha>0$. The proof of (5) is identical to the proof of [Tal92, Theorem $4.3.6$] by applying (4) and [Tal92, Theorem $2.2.23$], and hence omitted here.
Let $d_{2}^{\mathbb{U}}$ and $d_{\infty}^{\mathbb{U}}$ be the distances induced on ${\mathcal{F}}$ by the norms of $L_{2}$ and $L_{\infty}$ respectively, with respect to the measure induced by the uniform distribution on $[0,1]^{2}$. For all $f,g\in\mathcal{F}$,
$d_{2}^{\mathbb{D}}(f,g)$ $=\left(\int_{X}(f(x)-g(x))^{2}\mathrm{d}\mu_{\mathbb{D}}(x)\right)^{1/2}\leq\sqrt{\beta}\left(\int_{X}(f(x)-g(x))^{2}\mathrm{d}x\right)^{1/2}$
$=\sqrt{\beta}\left(\mathbb{E}_{x\sim\mathbb{U}}\left[(f(x)-g(x))^{2}\right]\right)^{1/2}=\sqrt{\beta}\cdot d_{2}^{\mathbb{U}}(f,g)\,,$
where the inequality holds by the assumption that $\mathbb{D}$ is $\beta$-smooth. Also, note that
$d_{\infty}^{\mathbb{D}}(f,g)\leq d_{\infty}^{\mathbb{U}}(f,g)=\sup_{x\in X}|f(x)-g(x)|\,.$
Then, by Lemma 14 and (5), together with the above two displayed inequalities, given that $\beta>0$ is a constant, we have
$\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\sup_{f\in{\mathcal{F}}}\left|Z_{f}(S)\right|\right]\leq O\left(\sqrt{N}\gamma_{2}\left({\mathcal{F}},d_{2}^{\mathbb{U}}\right)+\gamma_{1}\left({\mathcal{F}},d_{\infty}^{\mathbb{U}}\right)\right)\,.$
Recall that ${\mathcal{C}}$ denotes the class of $1$-lipschitz functions on $(X,\delta)$. Then, we obtain
$\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\sup_{f\in{\mathcal{C}}}\left|Z_{f}(S)\right|\right]\leq O\left(\sqrt{N}\gamma_{2}\left({\mathcal{C}},d_{2}^{\mathbb{U}}\right)+\gamma_{1}\left({\mathcal{C}},d_{\infty}^{\mathbb{U}}\right)\right)\leq O\left(\sqrt{n\log n}\right)\,,$ (6)
where the second inequality is implicitly shown in the proof of [Tal92, Theorem $4.3.2$]. Combining (3) and (6), it holds that
$\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\leq 2\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\sup_{f\in{\mathcal{C}}}\left|Z_{f}(S)\right|\right]\leq O\left(\sqrt{n\log n}\right)\,,$
which concludes the proof.
A.2 Proof of Lemma 3
@@
For $d=1$, let $F(x)=\int_{0}^{x}\mu_{\mathbb{D}}(x)\mathrm{d}x$ for $x\in[0,1]$ be the cumulative density function of $\mathbb{D}$. Let $S=\{s_{1},\ldots,s_{n}\}\sim\mathbb{D}^{n}$ and $R=\{r_{1},\ldots,r_{n}\}\sim\mathbb{D}^{n}$ be two sets of random points drawn from $\mathbb{D}$. Define $P_{S}=\{F(s_{1}),\ldots,F(s_{n})\}$ and $P_{R}=\{F(r_{1}),\ldots,F(r_{n})\}$, and each point in $P_{S}\cup P_{R}$ follows uniform distribution on $[0,1]$. Since the expected cost of the optimal matching between two sets of $n$ uniformly distributed points on $[0,1]$ is $\Theta(\sqrt{n})$, we have
$\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(P_{S},P_{R})\right]=\Theta\left(\sqrt{n}\right)\,.$ (7)
Note that for all $s,r\in[0,1]$,
@@
Let $M$ be the perfect matching between $S$ and $R$ with the minimum cost, and let $M^{\prime}$ be the perfect matching between $P_{S}$ and $P_{R}$ induced by $M$ such that $F(s)\in P_{S}$ is matched with $F(M(s))\in P_{R}$ for each $s\in S$. Hence, the cost of $M^{\prime}$ equals
@@
$\mathrm{OPT}(n)=\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\geq\frac{1}{\beta}\cdot\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(P_{S},P_{R})\right]\geq\Omega\left(\sqrt{n}\right)\,,$
@@
The lower bounds for $d\geq 2$ are proved by employing nearest-neighbor-distance, which is provided in the following lemma.
@@
Assume $\mathbb{D}$ to be $\beta$-smooth for some constant $\beta>0$. For all $d\geq 1$ and $S\subseteq[0,1]^{d}$ with $|S|=n$,
$\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]\geq\Omega\left(n^{-1/d}\right)\,.$
@@
Fix $S\subseteq[0,1]^{d}$ with $|S|=n$. For all $s\in S$ and $\ell>0$, let $B(s,\ell)=\{x\in[0,1]^{d}\mid\left\|{x-s}\right\|_{2}\leq\ell\}$ denote the ball restricted in $[0,1]^{d}$ with center $s$ and radius $\ell$, and we have
$\mathbb{P}_{r\sim\mathbb{D}}[\left\|{s-r}\right\|_{2}\leq\ell]=\int_{B(s,\ell)}\mu_{\mathbb{D}}(x)\mathrm{d}x\leq\int_{B(s,\ell)}\beta\mathrm{d}x\leq C\cdot\ell^{d}$
for some constant $C>0$. For all $s\in S$ and $r\in[0,1]^{d}$, let $O_{s,r}$ denote the event that $s\in\mathop{\arg\min}_{s^{\prime}\in S}\left\|{s^{\prime}-r}\right\|_{2}$ and $p_{s}=\mathbb{P}_{r\sim\mathbb{D}}[O_{s,r}]$. For all $s\in S$ and $0<\ell\leq(p_{s}/C)^{1/d}$, it holds that
@@
$\mathbb{E}_{r\sim\mathbb{D}}\left[\left\|{s-r}\right\|_{2}\mid O_{s,r}\right]$ $\geq\int_{0}^{(p_{s}/C)^{1/d}}\mathbb{P}_{r\sim\mathbb{D}}\left[\left\|{s-r}\right\|_{2}\geq\ell\mid O_{s,r}\right]\mathrm{d}\ell$
$=\frac{1}{p_{s}}\int_{0}^{(p_{s}/C)^{1/d}}\mathbb{P}_{r\sim\mathbb{D}}[O_{s,r}\land(\left\|{s-r}\right\|_{2}\geq\ell)]\mathrm{d}\ell$
$\geq\frac{1}{p_{s}}\int_{0}^{(p_{s}/C)^{1/d}}(p_{s}-C\cdot\ell^{d})\mathrm{d}\ell$
$=\frac{d}{d+1}\cdot\left(\frac{p_{s}}{C}\right)^{1/d}\,.$
@@
$\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]=\sum_{s\in S}\mathbb{P}_{r\sim\mathbb{D}}[O_{s,r}]\cdot\mathbb{E}_{r\sim\mathbb{D}}[\left\|{s-r}\right\|_{2}\mid O_{s,r}]\geq\sum_{s\in S}\Omega\left(p_{s}^{1+\frac{1}{d}}\right)\geq\Omega\left(n^{-1/d}\right)\,,$
where the last inequality holds since $\sum_{s\in S}p_{s}=1$ and $p_{s}\geq 0$ for every $s\in S$. ∎
Now, we are ready to prove the lower bounds for $d\geq 2$. By Lemma 15,
$\mathrm{OPT}(n)$ $=\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]=\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\mathbb{E}_{R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\right]$
$\geq\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[n\cdot\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]\right]\geq\Omega\left(n^{1-\frac{1}{d}}\right)\,,$
@@
A.3 Proof of Lemma 4
By Lemma 15 and the assumption that $|S|\leq O(n)$,
$\mathrm{OPT}(S,n)=\mathbb{E}_{R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\geq n\cdot\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]\geq\Omega\left(n^{1-\frac{1}{d}}\right)$
@@ B.1 Proof of Lemma 12
We use ${\mathcal{A}}$ to denote the greedy subroutine in the MNP algorithm. The tie-breaking rule of ${\mathcal{A}}$ is local by definition, and it remains to show that it is also non-increasing. Recall that $N_{{\mathcal{A}}}(r\mid S)$ is defined as the set of available servers closest to the request $r$ upon its arrival when the set of initial servers is $S$. For all sets of servers $S^{\prime},S^{\prime\prime}$ with $S^{\prime}\subseteq S^{\prime\prime}$, for every $s\in S^{\prime}$,
@@
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime}\right\}\geq\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}\right\}\,,$
which concludes that the tie-breaking rule of ${\mathcal{A}}$ is non-increasing.

### v3

@@ A.1 Proof of Lemma 2
For $d=1$, let $F(x):=\int_{0}^{x}\mu_{\mathbb{D}}(x)\mathrm{d}x$ for $x\in[0,1]$ be the cumulative density function of $\mathbb{D}$. Let $S=\{s_{1},\ldots,s_{n}\}\sim\mathbb{D}^{n}$ and $R=\{r_{1},\ldots,r_{n}\}\sim\mathbb{D}^{n}$ be two sets of random points independently drawn from $\mathbb{D}$. Define $P_{S}:=\{F(s_{1}),\ldots,F(s_{n})\}$ and $P_{R}:=\{F(r_{1}),\ldots,F(r_{n})\}$, and each point in $P_{S}\cup P_{R}$ follows the uniform distribution over $[0,1]$. Since the expected cost of the optimal matching between two sets of $n$ uniformly random points over $[0,1]$ is $\Theta(\sqrt{n})$, we have
$\mathbb{E}_{S\sim{\mathbb{D}}^{n},R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(P_{S},P_{R})\right]=\Theta\left(\sqrt{n}\right)\,.$ (1)
Note that for all $s,r\in[0,1]$, by the $\beta$-smoothness of $\mathbb{D}$},
@@
Let $M$ be the perfect matching between $S$ and $R$ with the minimum cost, and let $M^{\prime}$ be the perfect matching between $P_{S}$ and $P_{R}$ induced by $M$ such that $F(s)\in P_{S}$ is matched with $F(M(s))\in P_{R}$ for each $s\in S$. It follows that
@@
$\mathrm{OPT}_{\mathbb{D}}(n)=\mathbb{E}_{S\sim{\mathbb{D}}^{n},R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\geq\frac{1}{\beta}\cdot\mathbb{E}_{S\sim{\mathbb{D}}^{n},R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(P_{S},P_{R})\right]\geq\Omega\left(\sqrt{n}/\beta\right)\,$,
@@
The lower bound for $d\geq 2$ is proved by employing nearest-neighbor-distance, which is provided in the following lemma.
@@
Assume $\mathbb{D}$ to be a $\beta$-smooth distribution over $0,1]^{d}$ for $\beta\geq 1$. Then, for all $d\geq 1$ and $S\subseteq[0,1]^{d}$,
$\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]\geq\Omega\left(\beta^{-\frac{1}{d}}|S|^{-\frac{1}{d}}\right)\,.$
@@
Fix $S\subseteq[0,1]^{d}$. For all $s\in S$ and $\ell\geq 0$, let $B(s,\ell):=\{x\in[0,1]^{d}\mid\left\|{x-s}\right\|_{2}\leq\ell\}$ denote the ball restricted in $0,1]^{d}$ with center $s$ and radius $\ell$, and, by the $\beta$-smoothness of $\mathbb{D}$},
$\mathbb{P}_{r\sim\mathbb{D}}[\left\|{s-r}\right\|_{2}\leq\ell]=\int_{B(s,\ell)}\mu_{\mathbb{D}}(x)\mathrm{d}x\leq\int_{B(s,\ell)}\beta\mathrm{d}x\leq C\beta\cdot\ell^{d}$
for some constant $C>0$. For all $s\in S$ and $r\in[0,1]^{d}$, let $O_{s,r}$ denote the event that $s\in\mathop{\arg\min}_{s^{\prime}\in S}\left\|{s^{\prime}-r}\right\|_{2}$, and let $p_{s}:=\mathbb{P}_{r\sim\mathbb{D}}[O_{s,r}]$. For all $s\in S$ and $0<\ell\leq(p_{s}/C\beta)^{1/d}$, it holds that
@@
$\mathbb{E}_{r\sim\mathbb{D}}\left[\left\|{s-r}\right\|_{2}\mid O_{s,r}\right]$ $\geq\int_{0}^{(p_{s}/C\beta)^{1/d}}\mathbb{P}_{r\sim\mathbb{D}}\left[\left\|{s-r}\right\|_{2}\geq\ell\mid O_{s,r}\right]\mathrm{d}\ell$
$=\frac{1}{p_{s}}\int_{0}^{(p_{s}/C\beta)^{1/d}}\mathbb{P}_{r\sim\mathbb{D}}[O_{s,r}\land(\left\|{s-r}\right\|_{2}\geq\ell)]\mathrm{d}\ell$
$\geq\frac{1}{p_{s}}\int_{0}^{(p_{s}/C\beta)^{1/d}}(p_{s}-C\beta\cdot\ell^{d})\mathrm{d}\ell$
$=\frac{d}{d+1}\cdot\left(\frac{p_{s}}{C\beta}\right)^{1/d}\,.$
@@
$\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]$ $=\sum_{s\in S}\mathbb{P}_{r\sim\mathbb{D}}[O_{s,r}]\cdot\mathbb{E}_{r\sim\mathbb{D}}[\left\|{s-r}\right\|_{2}\mid O_{s,r}]$
$\geq\sum_{s\in S}\Omega\left(\beta^{-\frac{1}{d}}p_{s}^{1+\frac{1}{d}}\right)\geq\Omega\left(\beta^{-\frac{1}{d}}|S|^{-\frac{1}{d}}\right)\,$,
where the last inequality holds by the convexity of $f(x)=x^{1+\frac{1}{d}}$ for $x\geq 0$. ∎
Now, we are ready to prove the lower bound for $d\geq 2$. By Lemma 14,
$\mathrm{OPT}_{\mathbb{D}}(n)$ $=\mathbb{E}_{S\sim{\mathbb{D}}^{n},R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]=\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\mathbb{E}_{R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\right]$
$\geq\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[n\cdot\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]\right]\geq\Omega\left(\beta^{-\frac{1}{d}}n^{1-\frac{1}{d}}\right)\,$,
@@
A.2 Proof of Lemma 3
By Lemma 14 and the assumption that $|S|=\kappa n$,
$\mathrm{OPT}_{\mathbb{D}}(S,n)=\mathbb{E}_{R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\geq n\cdot\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]\geq\Omega\left((\beta\kappa)^{-\frac{1}{d}}n^{1-\frac{1}{d}}\right)\,$,
@@ B.1 Proof of Lemma 12
We use $\mathcal{A}$} to denote the greedy subroutine in the MNP algorithm. The tie-breaking rule of $\mathcal{A}$} is local by definition, and it remains to show that it is non-increasing. Recall that $N_{{\mathcal{A}}}(r\mid S)$ denotes the set of available servers closest to the request $r$ upon its arrival when the set of initial servers is $S$. For all sets of servers $S^{\prime},S^{\prime\prime}$ with $S^{\prime}\subseteq S^{\prime\prime}$, for every $s\in S^{\prime}$,
@@
[…]
implying that the tie-breaking rule of $\mathcal{A}$} is non-increasing.